Given the following two fuel samples, please calculate (show all work) the fuel moisture content of each sample and answer the following questions (Tare weight = 10.0 gm)

Sample 1: Wet weight = 76.6 g, Dry weight = 70.8 g

Sample 2: Wet weight = 110.5 g, Dry weight = 51.2 g.

a. Which sample is the driest?

b. How would you classify Sample 1 and Sample 2 (fuel type)?

c. Which fuel would be easier to ignite and why?

d. What type of vegetation (fuel) would you expect Sample 1 to be? Sample 2?
Answer #1

Thank you for asking. Let's answer each questions one by one in the best way possible.

We are given -

Tare weight= 10 gram

Sample 1 - wet weight = 76.6 gram

dry weight = 70.8 gram

Fuel moisture content of sample 1 = 76.6-70.8/70.8-10*100 = 5.8/ 60.8*100 = 8.22%

Sample 2 - wet weight = 110.5 gram

dry weight = 51.2 gram

Fuel moisture content of sample 2 = 110.5-51.2/51.2-10*100 = 59.3/41.2*100= 143.93%

A.) From the above calculation, its clear that Sample 1 is the driest as it has moisture content of mere 8.22 %.

B.) Sample 1 can be classified as dead fuel moisture while Sample 2 can be classified as live fuel mositure.

C.) Sample 1 fuel will be easier to ignite as compared to sample 2 because of less moisture content. Because of the low moisture content, heat energy can be directly utilized to burn the flames unlike in the fuels having high moisture content where much of the heat energy gets utilized in the process of evaporation.

D.) Sample 1 can include vegetations like dead grasses, frost killed leaves etc and are dead fuel moisture while Sample 2 fuel includes vegetation like conifers, shrubs, green grasses etc., and are live fuel moisture.

I hope this will help you. In case if you have any problem, please feel free to ask. I'll be more than happy to help. Thanks again.